# Gravitational waves of quasi-circular, inspiraling black hole binaries in an ultralight vector dark-matter environment

### v1

@@ IV Impact of Ultralight Dark Matter Vector field on the Orbital Motion of a Compact Binary
The contribution from the binary system to the linearized Riemann tensor in Eq. (16), ${R^{i}}_{0j0}[h^{M}_{\mu\nu}]$, yields the post-Newtonian equations of motion in the center of mass frame. In particular, when the metric perturbation $h_{\mu\nu}^{M}$ is taken to leading post-Newtonian order, one obtains the usual Newtonian law of gravity, i.e. ${R^{i}}_{0j0}[h^{M}_{\mu\nu}]r^{j}=-({G\,M}/{r^{2}})r^{i}$ (see for instance Sec. 17.4 in [36]). The contribution from the vector field can then be interpreted as a perturbation to the post-Newtonian equations of motion. The components of ${R^{i}}_{0j0}[h^{A}_{\mu\nu}]$ can be written in terms of the vector field fluid variables using Eq. (15), as
@@
where we recall that ${\hat{X}^{i}}_{\,\,j}={\delta^{i}}_{j}-2\hat{A}^{i}\hat{A}_{j}$. Moreover, since the unperturbed problem reduces to Newtonian gravity, we know that Kepler’s laws are valid to leading order, so the frequency of the orbit is $\omega^{2}\sim M/r^{3}+{\cal{O}}(\rho_{A})$, and the binding energy of the orbit is the usual Newtonian potential, namely $E_{M}\sim-GM/2r+{\cal{O}}(\rho_{A})$.
@@
where $\hat{z}$ is normal to the orbital plane. In this basis, one can show that the equations of motion for the semi-major axis of a circular orbit is given by [^5] For a detailed derivation of the osculating equations, see for instance Sec. 3.3 of Ref. [37].
@@
The equation of motion for the semi-major axis (Eq. Eq. 23) deserves further discussion. First, we see that, for circular orbits, we have a non-zero perturbation of the orbit, as long as the perturbing forces have tangential components. Scalar field models do not have tangential components due to the symmetry of the field. Fields of higher spin, however, can break this symmetry, thus producing a non-vanishing effect, even for circular orbits. Note also that the perturbing force is conservative, so the temporal evolution of the semi-major axis is also conservative and, in particular, it is not driven by dissipative forces. Gravitational-wave back-reaction is a dissipative effect that also induces a time-variability in the semi-major axis, forcing the binary to inspiral, but this back-reaction effect is qualitatively different from the conservative effect of the perturbing vectorial force studied here.
@@
How large is this perturbation relative to other perturbations we expect the binary system to experience? The main other perturbation is gravitational-wave backreaction, which induces a change in the semi-major axis of the form [38]
$\left(\frac{\dot{b}}{b}\right)_{GW}=-\frac{64}{5}\frac{G^{3}\eta}{c^{5}b}\left(\frac{M}{b}\right)^{3}\,,$ (26)
@@
$\frac{\left({\dot{b}}/{b}\right)_{A}}{\left({\dot{b}}/{b}\right)_{GW}}\sim\frac{c^{3}b^{4}}{G^{2}\eta M^{3}\omega}\rho_{A}\,.$ (27)
From this ratio, we observe that the effects of the vector field are enhanced for asymmetric binaries that are widely separated, as we anticipated in the introduction.
@@ V.1 The stationary-phase approximation
[…]
@@
where we used the balance law $\dot{E}\equiv\mathcal{L}$ to relate the rate of change of the orbital energy of the binary $\dot{E}$ to the gravitational-wave luminosity $\mathcal{L}$. With this in hand, the Fourier phase is given by
[…]
@@
where $E_{M}$ and $\mathcal{L}_{M}$ are the binding energy of the orbit and luminosity of the gravitational wave respectively in general relativity and to as high a post-Newtonian order as one wishes to include. The quantities $\delta E$ and $\delta\mathcal{L}$ represent corrections to $E$ and ${\cal{L}}$ that are not from the binary. In this work, we focus on corrections that are only due to the presence of a vector-field dark-matter environment. As we explained earlier, the vector-field effect is conservative, because the perturbing force can be derived from the spatial gradient of a potential that is even under time-reversal. In particular, the vector-field environment that we consider does not induce radiative effects, since on the scales of interest to use, the vector-field is not a propagating wave that carries energy-momentum away from the binary. Therefore, the energy carried away in gravitational waves is the only source of energy loss. With this in mind, we conclude that $\delta\mathcal{L}=0$ and $\delta E=V_{A}$, where $V_{A}$ is given in Eq. (21).
@@
where $\Psi_{M}(f)$ is the Fourier gravitational-wave phase to as high a post-Newtonian order as one wishes to include, while $\Psi_{A}(f)$ is the dephasing correction generated by the vector-field environment. To find $\Psi_{A}(f)$, we insert the perturbed binding energy $V_{A}$ of Eq. (41) into Eq. (41) to obtain
[…]
@@
There is another simple way to calculate the dephasing which consists in calculating the perturbation in the orbital frequency of the binary as a function of time due to the presence of the vector field. That is, we split the variation of the frequency in a contribution from general relativity plus the perturbation from the vector field, $\dot{F}=\dot{F}_{M}+\dot{F}_{A}$, with $\dot{F}_{M}\gg\dot{F}_{A}$. Then, we can calculate the dephasing as
[…]
We can now use Kepler law to relate the variation in the frequency with a variation in the semi-major axes, which we calculated in Eq. (23), as $\dot{F}_{A}=\tfrac{3}{2}F(\dot{b}/{b})_{A}$. In this way, we obtain an equivalent expression as the one given in Eq. (45).
@@ V.2 Calculation of the dephasing $\Psi_{A}(f)$
In order to solve the integral of Eq. (45), we need to calculate the variation of the binding energy with respect to the frequency. To do so, we follow the method of osculating orbits, and assume that the binding energy, to leading order, is given by the Newtonian expression, $E_{M}=-GM/2b$. Then, the variations in the binding energy of the system are given by variations in the semi-major axes. Using the chain rule, we have that
$\frac{dE_{A}}{dF}=\left(\frac{dE_{A}}{dt}\right)\left(\frac{dt}{dF}\right)=-E_{M}\left(\frac{dt}{dF}\right)_{GW}\left(\frac{\dot{b}}{b}\right)_{A}\,,$ (48)
@@
$\left(\frac{dt}{dF}\right)_{GW}=-\frac{dE_{M}}{dF}\frac{1}{{\cal{L}}_{GW}}=-\frac{5\pi}{48}\,\eta^{-1}M^{2}(2\pi MF)^{-\frac{11}{3}}\,.$ (49)
Inserting the above expressions in the integral of Eq. (45), and changing variables to $x=(2\pi F\mathcal{M})^{\frac{1}{3}}$, we have
@@
×$\times\frac{1}{x^{20}}\cos\left(\frac{5\mathcal{M}m_{A}}{128x^{8}}+\gamma\right)\sin\left(\frac{5}{128x^{5}}-2\varphi\right)\,,$ (50)
where we typically have $0.30.01<x<$ 0.3 for systems in the LISA band.
@@
We look for stationary points where $\phi^{\prime}(x_{\rm{stp}})=0$, which yields $\pi f_{\rm{stp}}=8m_{A}/5$. Observe that the stationary point depends on the vector field mass $m_{A}$, but it is independent of the total mass of the binary $M$. Since we are integrating from $F=+\infty$, the integral is approximately zero until the stationary point $F\sim f_{\rm{stp}}$ is reached, so $\delta\psi(f)=0$ for $F\gg f_{\rm{stp}}$. Once the stationary point is crossed, the integral has a contribution from the neighborhood of $f_{\rm{stp}}$ given by
@@
×$\times\sin(\alpha)^{2}\sin(\xi-\frac{\pi}{4}+2\varphi)\,,$
@@
Figure 3: Dephasing of the gravitational wave calculated as in Eq. (50) under the stationary-phase approximation (black dashed line) and numerically integrated (red dots), for $m=10^{-18}\rm{eV}$, $M=10^{5}\rm{\rm{M}_{\odot}}$ and $\eta=0.25$. The vertical dash-dot gray line corresponds to the stationary point $\pi f_{\rm{stp}}=\frac{8}{5}m_{A}$. For $f>f_{\rm{stp}}$ the integral averages to zero. Observe that the stationary-phase approximation is very accurate, with percentage errors $\ll 1\%$ in the range shown.
@@ VI Fisher analysis
[…]
where $S_{n}(f)$ is the power noise spectrum of the instrument, the overhead tilde stands for the Fourier transform, and the superscript star for complex conjugation. In this work, we focus on binary systems that can be detected in the LISA band, so we perform our analyses using an approximation to the LISA power noise spectrum [39].
@@
[…]
@@
$\Psi(f)=\beta\left(1-\frac{f}{f_{\rm{stp}}}\right)\theta(f_{\rm{stp}}-f)\,,$ (58)
@@
We then add these two parameters to the rest of the parameters contained in the vacuum part of the gravitational wave signal $\tilde{h}_{M}$ in Eq. (47), which in this paper we will model through a TaylorF2 model [40] valid only up to the ISCO. Equation (47) then gives us our dark-matter enhanced waveform model.
@@
×$\qquad\times\delta(f-f_{\rm{stp}})\bigg{]}\,\tilde{h}(f)\,.$
@@
Figure 4 shows the gravitational-wave characteristic strain for different binary masses as a function of the gravitational-wave frequency. The black solid curve is the characteristic LISA sensitivity strain [39]. For each binary mass, we show the frequency range each binary sweeps through. The initial frequency $f_{\rm{min}}$ is calculated as the frequency the binary had four year before ISCO, corresponding to a large observation period with LISA. Observe that each frequency range also corresponds to the mass range of the vector field that could be tested with a gravitational-wave observation. This is because, as mentioned before, for each binary evolution, the orbital frequency resonates with the vector field mass at some point in the evolution of the binary, activating the step function and thus breaking the degeneracy. In particular, we see that for systems that could be probed by LISA, the mass range for the vector field that could be tested in a detection is $m_{A}\sim(10^{-19},10^{-15})\rm{eV}$.
Figure 4: LISA’s sensitivity strain (black curve). The different colors indicate the gravitational wave characteristic strain for each binary mass during the inspiral. The dashed lines indicate $f_{\rm{min}}$ for each system, calculated as the frequency four years before the ISCO is reached (corresponding to LISA’s observation period). We assumed that the systems where at $D_{L}=1\rm{Gpc}$. On the top axes we show $\omega_{GW}=2\pi f_{GW}$.
@@ VII Conclusions
We performed a Fisher analysis to estimate the accuracy to which we may be able to constrain the dark-matter density $\rho_{A}$ with a future LISA observation. We found that 4-year LISA observations should allow us to measure dark-matter densities larger than $10^{16}\;{\rm{GeV}}/{\rm{cm}}^{3}$ for vector masses in the range $(10^{-19},10^{-16})\;{\rm{eV}}$. We also find that these measurements are best for asymmetric binaries with low total mass. Interestingly, some of these are systems whose merger could potentially also be observed by LIGO, after LISA observes the early inspiral [41]. Such multi-wavelength gravitational-wave events seem ideal to probe non-vacuum or beyond-Einstein effects in gravitational wave signals [42, 43]. Therefore, future LISA observations of the gravitational waves emitted in the quasi-circular inspiral of massive black holes may detect or yield interesting constraints on the density of vector-type ultralight dark matter surrounding binary black holes systems.
Future work could focus on a complete study of the perturbations in the gravitational waveform due to the presence of a vector field environment. Such a study should include perturbations in the waveform resulting from the variation of other orbital parameters, in addition to the variation of the semi-major axes considered here. In particular, future work could focus on the variation of the inclination of the orbital plane, which has an impact on the amplitude of the gravitational wave. Also, the formalism presented in this manuscript can be generalized for eccentric orbits [44], since the vector field also has an impact on the evolution of the eccentricity of the orbits.

### v2

@@ IV Impact of Ultralight Dark Matter Vector field on the Orbital Motion of a Compact Binary
The contribution from the binary system to the linearized Riemann tensor in Eq. (18), ${R^{i}}_{0j0}[h^{M}_{\mu\nu}]$, yields the post-Newtonian equations of motion in the center of mass frame. In particular, when the metric perturbation $h_{\mu\nu}^{M}$ is taken to leading post-Newtonian order, one obtains the usual Newtonian law of gravity, i.e. ${R^{i}}_{0j0}[h^{M}_{\mu\nu}]r^{j}=-({G\,M}/{r^{2}})r^{i}$ (see for instance Sec. 17.4 in [Misner:1973prb]). The contribution from the vector field can then be interpreted as a perturbation to the post-Newtonian equations of motion. The components of ${R^{i}}_{0j0}[h^{A}_{\mu\nu}]$ can be written in terms of the vector field fluid variables using Eq. (17), as
@@
where we recall that ${\hat{X}^{i}}_{\,\,j}={\delta^{i}}_{j}-2\hat{A}^{i}\hat{A}_{j}$. Moreover, since the unperturbed problem reduces to Newtonian gravity, we know that Kepler’s laws are valid to leading order, so the frequency of the orbit is $\omega^{2}\sim M/r^{3}+{\cal{O}}(\rho_{A})$, and the binding energy of the orbit is the usual Newtonian potential, namely $E_{M}\sim-G\mu M/2r+{\cal{O}}(\rho_{A})$.
@@
where $\hat{z}$ is normal to the orbital plane. In this basis, one can show that the equations of motion for the semi-major axis of a circular orbit is given by [^5] For a detailed derivation of the osculating equations, see for instance Sec. 3.3 of Ref. [Poisson_Will_2014].
@@
The equation of motion for the semi-major axis (Eq. 25) deserves further discussion. First, we see that, for circular orbits, we have a non-zero perturbation of the orbit, as long as the perturbing forces have tangential components. Scalar field models do not have tangential components due to the symmetry of the field. Fields of higher spin, however, can break this symmetry, thus producing a non-vanishing effect, even for circular orbits. Note also that the perturbing force is conservative, so the temporal evolution of the semi-major axis is also conservative and, in particular, it is not driven by dissipative forces. Gravitational-wave back-reaction is a dissipative effect that also induces a time-variability in the semi-major axis, forcing the binary to inspiral, but this back-reaction effect is qualitatively different from the conservative effect of the perturbing vectorial force studied here.
@@
How large is this perturbation relative to other perturbations we expect the binary system to experience? The main other perturbation is gravitational-wave backreaction, which induces a change in the semi-major axis of the form [Peters:1964zz]
$\left(\frac{\dot{b}}{b}\right)_{M}=-\frac{64}{5}\frac{G^{3}\eta}{c^{5}b}\left(\frac{M}{b}\right)^{3}\,$, (28)
@@
$\frac{\left({\dot{b}}/{b}\right)_{A}}{\left({\dot{b}}/{b}\right)_{M}}\sim\frac{c^{3}b^{4}}{G^{2}\eta M^{3}\omega}\rho_{A}$ (29)
$\quad\sim 1.5\times 10^{-16}\frac{\eta^{-1}\rho_{A}}{0.3\,\frac{\mathrm{GeV}}{\mathrm{cm^{3}}}}\left(\frac{10^{6}\mathrm{\rm{M}_{\odot}}}{M}\right)^{5/3}\left(\frac{10^{-5}\mathrm{Hz}}{f}\right)^{8/3}\,$,
where we used the solar system dark matter energy density as the reference value for $\rho_{A}$. From this ratio, we observe that the effects of the vector field are enhanced for asymmetric binaries that are widely separated, as we anticipated in the introduction.
@@ V.1 The stationary-phase approximation
$\Phi(t_{*})$ $=\Phi_{c}-2\pi\int^{\infty}_{f/2}\frac{F(F^{\prime})}{\dot{F}(F^{\prime})}dF^{\prime}\,$, (39)
$t_{*}$ $=t_{c}-\int^{\infty}_{f/2}\frac{1}{\dot{F}(F^{\prime})}dF^{\prime},$ (40)
@@
where we used the balance law $\dot{E}\equiv$-$\mathcal{L}$ to relate the rate of change of the orbital energy of the binary $\dot{E}$ to the gravitational-wave luminosity $\mathcal{L}$. With this in hand, the Fourier phase is given by
$\Psi(f)=2\pi ft_{c}-2\Phi_{c}-2\pi\int^{\frac{f}{2}}_{\infty}(2F-f)\mathcal{L}^{-1}\frac{dE}{dF}\,dF\,.$ (43)
@@
[…]
@@
where $\Psi_{M}(f)$ is the Fourier gravitational-wave phase to as high a post-Newtonian order as one wishes to include, while $\Psi_{A}(f)$ is the dephasing correction generated by the vector-field environment. To find $\Psi_{A}(f)$, we insert the perturbed binding energy $\delta E=V_{A}$ of Eq. (23) into Eq. (43) to obtain
$\Psi_{A}(f)=-2\pi\int^{\frac{f}{2}}(2F-f)\mathcal{L}_{\rm{M}}^{-1}\frac{dV_{A}}{dF}\,dF\,.$ (47)
@@
There is another simple way to calculate the dephasing which consists in calculating the perturbation in the orbital frequency of the binary as a function of time due to the presence of the vector field. That is, we split the variation of the frequency in a contribution from general relativity [^6] At leading order in post newtonian perturbation theory, the variation of the frequency in general relativity is given by $\dot{F}_{M}=\tfrac{3}{2}F(\dot{b}/{b})_{M}$, with $(\dot{b}/{b})_{M}$ given in Eq. (28). plus the perturbation from the vector field, $\dot{F}=\dot{F}_{M}+\dot{F}_{A}$, with $\dot{F}_{M}\gg\dot{F}_{A}$. Then, we can calculate the dephasing as
$\Psi_{A}(f)=-2\pi\int^{\frac{f}{2}}(2F-f)\frac{\dot{F}_{A}}{\dot{F}_{M}^{2}}dF\,.$ (48)
We can now use Kepler law to relate the variation in the frequency with a variation in the semi-major axes, which we calculated in Eq. (25), as $\dot{F}_{A}=-\tfrac{3}{2}F(\dot{b}/{b})_{A}$. In this way, we obtain an equivalent expression as the one given in Eq. (47).
@@ V.2 Calculation of the dephasing $\Psi_{A}(f)$
In order to solve the integral of Eq. (47), we need to calculate the variation of the binding energy with respect to the frequency. To do so, we follow the method of osculating orbits, and assume that the binding energy, to leading order, is given by the Newtonian expression, $E_{M}=-G\mu M/2b$. Then, the variations in the binding energy of the system are given by variations in the semi-major axes. Using the chain rule, we have that
$\frac{dE_{A}}{dF}=\left(\frac{dE_{A}}{dt}\right)\left(\frac{dt}{dF}\right)=-E_{M}\left(\frac{dt}{dF}\right)_{M}\left(\frac{\dot{b}}{b}\right)_{A}\,$, (50)
@@
$\left(\frac{dt}{dF}\right)_{M}=-\frac{dE_{M}}{dF}\frac{1}{{\cal{L}}_{M}}=-\frac{5\pi}{48}\,\eta^{-1}M^{2}(2\pi MF)^{-\frac{11}{3}}\,.$ (51)
Inserting the above expressions in the integral of Eq. (47), and changing variables to $x=(2\pi F\mathcal{M})^{\frac{1}{3}}$ (with $\mathcal{M}=\eta^{3/5}M$ the chirp mass), we have
@@
$\times\frac{1}{x^{20}}\cos\left(\frac{5\mathcal{M}m_{A}}{128x^{8}}+\gamma\right)\sin\left(\frac{5}{128x^{5}}-2\varphi\right)\,$, (52)
where we typically have $0.01<x<0.3$ for systems in the LISA band.
@@
We look for stationary points where $\phi^{\prime}(x_{\rm{stp}})=0$, which yields $\pi f_{\rm{stp}}=8m_{A}/5$. Observe that the stationary point depends on the vector field mass $m_{A}$, but it is independent of the total mass of the binary $M$. Since we are integrating from $F=+\infty$, the integral is approximately zero until the stationary point $F\sim f_{\rm{stp}}$ is reached, so $\Psi_{A}(f)=0$ for $F\gg f_{\rm{stp}}$. Once the stationary point is crossed, the integral has a contribution from the neighborhood of $f_{\rm{stp}}$ given by
@@
$\times\sin(\alpha)^{2}\sin(\xi-\frac{\pi}{4}+2\varphi)\,$,
@@
Figure 3: Dephasing of the gravitational wave calculated as in Eq. (52) under the stationary-phase approximation (black dashed line) and numerically integrated (red dots), for $m=10^{-18}\rm{eV}$, $M=10^{5}\rm{\rm{M}_{\odot}}$, $\eta=0.25$, $\varphi=0$ and $\alpha=\pi/4$. The vertical dash-dot gray line corresponds to the stationary point $\pi f_{\rm{stp}}=\frac{8}{5}m_{A}$. For $f>f_{\rm{stp}}$ the integral averages to zero. Observe that the stationary-phase approximation is very accurate, with percentage errors $\ll 1\%$ in the range shown.
@@ VI Fisher analysis
$\langle h|g\rangle=2\int_{f_{\rm{min}}}^{f_{\rm{max}}}\frac{\tilde{h}^{*}\tilde{g}+\tilde{h}\tilde{g}^{*}}{S_{n}}df\,$, (56)
where $S_{n}(f)$ is the power noise spectrum of the instrument, the overhead tilde stands for the Fourier transform, and the superscript star for complex conjugation. In this work, we focus on binary systems that can be detected in the LISA band, so we perform our analyses using an approximation to the LISA power noise spectrum [Robson:2018ifk].
@@
where $\Sigma^{aa}$ is the $(a,a$) diagonal element of the inverse of the covariance matrix (and summation is not implied here). The covariance matrix, also known as the Fisher information matrix $\Gamma_{ab}$, is defined as
@@
$\Psi_{A}(f)=\beta\left(1-\frac{f}{f_{\rm{stp}}}\right)\theta(f_{\rm{stp}}-f)\,$, (60)
@@
We then add these two parameters to the rest of the parameters contained in the vacuum part of the gravitational wave signal $\tilde{h}_{M}$ in Eq. (49), which in this paper we will model through a TaylorF2 model [Damour:2000zb] valid only up to the ISCO, namely
$\Psi_{M}$ $=2\pi ft_{c}-\Phi_{c}-\frac{\pi}{4}+\frac{3}{128}U^{-5/3}$ (62)
$-\frac{3\pi}{8}\eta^{-3/5}U^{-2/3}+\frac{5}{96}\eta^{-2/5}\left(\frac{743}{336}+\frac{11\eta}{4}\right)U^{-1}\,$,
with $U=\pi\mathcal{M}f$. Equation (49) then gives us our dark-matter enhanced waveform model.
@@
$\qquad\times\delta(f-f_{\rm{stp}})\bigg]\,\tilde{h}(f)\,.$
@@
Figure 4 shows the gravitational-wave characteristic strain for different binary masses as a function of the gravitational-wave frequency. The black solid curve is the characteristic LISA sensitivity strain [Robson:2018ifk]. For each binary mass, we show the frequency range each binary sweeps through. The initial frequency $f_{\rm{min}}$ is calculated as the frequency the binary had four year before ISCO, corresponding to a large observation period with LISA. Observe that each frequency range also corresponds to the mass range of the vector field that could be tested with a gravitational-wave observation. This is because, as mentioned before, for each binary evolution, the orbital frequency resonates with the vector field mass at some point in the evolution of the binary, activating the step function and thus breaking the degeneracy. In particular, we see that for systems that could be probed by LISA, the mass range for the vector field that could be tested in a detection is $m_{A}\sim(10^{-19},10^{-15})\rm{eV}$.
Figure 4: LISA’s sensitivity strain (black curve). The different colors indicate the gravitational wave characteristic strain for each binary mass during the inspiral, defined as $|h_{c}(f)|=2f|\mathcal{A}(f)|$. We consider symmetric-mass binaries at $D_{L}=1\rm{Gpc}$. The mass of each member of the binary is the one indicated in the label. The dashed lines indicate $f_{\rm{min}}$ for each system, calculated as the frequency four years before the ISCO is reached (corresponding to LISA’s observation period). On the top axes we show $\omega_{GW}\equiv 2\pi f$.
@@ VII Conclusions
We performed a Fisher analysis to estimate the accuracy to which we may be able to constrain the dark-matter density $\rho_{A}$ with a future LISA observation. We found that 4-year LISA observations should allow us to measure dark-matter densities larger than $10^{16}\;{\rm{GeV}}/{\rm{cm}}^{3}$ for vector masses in the range $(10^{-19},10^{-16})\;{\rm{eV}}$. We also find that these measurements are best for asymmetric binaries with low total mass. Interestingly, some of these are systems whose merger could potentially also be observed by LIGO, after LISA observes the early inspiral [Sesana:2016ljz]. Such multi-wavelength gravitational-wave events seem ideal to probe non-vacuum or beyond-Einstein effects in gravitational wave signals [Barausse:2016eii, Perkins:2020tra]. Therefore, future LISA observations of the gravitational waves emitted in the quasi-circular inspiral of massive black holes may detect or yield interesting constraints on the density of vector-type ultralight dark matter surrounding binary black holes systems.
Future work could focus on a complete study of the perturbations in the gravitational waveform due to the presence of a vector field environment. Such a study should include perturbations in the waveform resulting from the variation of other orbital parameters, in addition to the variation of the semi-major axes considered here. In particular, future work could focus on the variation of the inclination of the orbital plane, which has an impact on the amplitude of the gravitational wave. Also, the formalism presented in this manuscript can be generalized for eccentric orbits [Yunes:2009yz], since the vector field also has an impact on the evolution of the eccentricity of the orbits.